) The lattice energy of magnesium bromide, MgBr2, can be calculated using the enthalpy changes shown in the table. Type of enthalpy change Value of enthalpy change/ kJ molˉ1 1 st ionization energy of magnesium (IE1) +736 2 nd ionization energy of magnesium (IE2) +1450 1 st electron affinity of bromine (EA) –325 Enthalpy change of formation of MgBr2 (∆Hf) –524 Enthalpy change of atomization of magnesium (∆Hat) +150 Enthalpy change of atomization of bromine (∆Hat) +112 (i) State the meaning of the terms; –Lattice energy; –2 nd ionization energy; –1 st electron affinity.


1
Expert's answer
2021-01-19T04:40:45-0500

The meaning of the terms:

–Lattice energy is energy (equals to enthalpy at constant pressure) required for the destruction of 1 mole ionic solid lattice producing gas-phase ions (cations and anions):

AxBy(s) = x An+(g) + y Bm-(g);

–2 nd ionization energy is the minimum energy required for withdrawal of one electron from a gas-phase single-charged cation:

A+(g) = A2+(g) + e-;

–1 st electron affinity is energy change after acception of one electron by a gas-phase atom:

B(g) + e- = B-(g)


In this example, lattice energy is calculated as:

U = 736+1450-2*325+524+150+112 = 2322 kJ/mol